OpenStudy (anonymous):

What 5-digit number has the following features: If we put the numeral 1 at the beginning, we get a number three times smaller than if we put the numeral 1 at the end of the number.

OpenStudy (anonymous):

1**** = 10000 + **** ####1 = ####0 + 1 10000 + **** = 3(####) + 3 9997 = 3(####) - **** lol i dunno why i used such a notation

OpenStudy (anonymous):

that looks like a 4 digit number

OpenStudy (anonymous):

its first digit is 3??

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

6 or 9

OpenStudy (anonymous):

there's a simple closed form expression that solves it

OpenStudy (anonymous):

i give up:(

OpenStudy (anonymous):

Using an easy equation: 3(100000 + x) = 10x+1 (Why? Well, adding 100000 puts a 1 at the front of a five-digit number, and multiplying by 10 and adding 1 puts a 1 at the end of a number) Solving this gives: 10x+1 = 3(100000 + x) 10x+1 = 300000 + 3x 10x = 299999 + 3x 7x = 299999 x= 299999/7 = 42857 The answer is 42857.

OpenStudy (anonymous):

why did you post it ????? -___-

OpenStudy (anonymous):

no one else seeemd to be working on it

OpenStudy (anonymous):

i was getting the concept :\

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

lol it's alright :-)